OpenStudy (anonymous):

Help please. Evaluate the integral by making the substitution. The integral of (x+2)/(x^2+4x) dx u=x^2+4x

OpenStudy (zepp):

You can replace the x^2+4x in your integral for u, and derive u, du=2x+4dx; 2du=x+2dx

OpenStudy (zepp):

Then you can proceed to solve for your integral: \[\int \frac{2}{u}du\]

OpenStudy (dan815):